Circular No. 6592

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

MARKARIAN 501 A. C. Breslin, P. Boyle, J. Buckley, M. Catanese, F. Krennrich, J.

McEnery, J. Quinn, J. Rose, R. Srinivasan, and T. C. Weekes report for the Whipple Observatory Gamma Ray Collaboration: “TeV emission from the BL Lac object Mkn 501 has been in a high state since Feb. 14, when its flux was observed to be comparable to that from the Crab nebula. Thirty hours of observations were made with the Whipple Observatory 10-m reflector (+ 151-pixel camera) at an energy threshold of 300 GeV. Since Mar. 5, the emission level has increased, reaching a peak on Mar. 13 at 2–2.5 times that of the Crab in a 4-hr observation. By observing over a range of elevations, we observed a broad spectrum of energies; the spectrum is steeper than that of the Crab and Mkn 421 above 2 TeV, and it extends at least to energies of 5 TeV (5σ significance). The flux is variable on a timescale of hours to days, but on average from Feb. 10 through Mar. 15 is comparable to the Crab at the threshold energy; at discovery in 1995, the flux was 0.08 times that of the Crab (Quinn et al. 1996, Ap.J. 456, L83).”

B. Degrange for the CAT group [Čerenkov Array at Themis, incorporating the Centre d’Etudes Nucleaire de Bordeaux-Gradignan; Centre d’Etude Spatiale des Rayonnements, Toulouse; Laboratoire de Physique Corpusculaire, College de France; Laboratoire de Physique Nucleaire de Hautes Energies at Ecole Polytechnique and Paris; Service d’Astrophysique, Centre d’Etudes de Saclay; and Universities of Perpignan, Charles, and Olmouc], operating a very-high-energy (VHE) γ-ray detector in the French Pyrenees, have observed Mkn 501 in a high state for 18.55 hr (and a control region for 4.54 hr) on Mar. 7–13 and 17–18. The maximum of emission was observed on Mar. 13, with a flux at a level of 4 times that of the Crab Nebula, for which the emission was seen over the energy range from 0.3 to > 5 TeV; on this particular night, the emission seemed stable at this high level during the observation time (1.71 hr). Also, variations on a half-hour timescale were seen on Mar. 10.

D. Kranich, E. Lorenz, D. Petry, and G. Rauterberg report that the High Energy Gamma Ray Array (HEGRA) collaboration, operating a cosmic-ray detector complex for VHE and ultra-high-energy γ-ray astronomy on La Palma, has observed unusually strong VHE γ-ray emission from Mkn 501 with a 5-m^{2} air Čerenkov telescope on every night during Mar. 10–17. A mean flux of ∼ 1.5 × 10^{−11} cm^{−2} s^{−1} was observed above 2.5 TeV. The flux is ∼ 10 times higher than observed in a 200-hr observation in 1996. The spectrum extends above 7 TeV. Observations are continuing with the full HEGRA array of six telescopes.

1997 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green